I have to solve this equation : p + 9 = 2
May I please have some help?

Found 2 solutions by THANApHD, praseenakos@yahoo.com:
Answer by THANApHD(104)


p+9 = 2
u hav to solve for "p". So u got to hav "p" on one side of equal sign and the other side with the numbers
our only obstacle here is 9, which shares the space with "p". So eliminate it.
p+9-9 = 2-9 (subtract 9 from both sides)
p = -7

tada.... u got the answer

Answer by praseenakos@yahoo.com(507)

Question:

solve this equation : p + 9 = 2

Answer:

p + 9 = 2

Subtract 9 from both sides of the expression...


==> p + 9 - 9 = 2 - 9


==> p = -7


Hence the solution.

Hope you understood.



Regards.
Praseena.